Crochet Company allocates overhead based on machine hours. Estimated overhead costs for the year total $390,000. The company estimates that it will use 60,000 machine hours during the year, have $500,000 in materials costs, and $400,000 in direct labor costs. Crochet actually has $380,000 of overhead, works 58,000 machine hours, has $550,000 in materials costs, and $420,000 in direct labor costs.

REQUIRED: a. What is the overhead application rate for the year? b. What is the amount of applied overhead for the year? c. What is the amount of under- or over-applied overhead for the year? d. Prepare the journal entry to record the under- or over-applied overhead.
Answer #1

a. Overhead application rate = 390,000 / 60,000 = $6.50 per machine hour.



b. Applied overhead = 6.50 x 58,000 = $377,000


c. Applied overhead is less than actual overhead, so overhead is under-applied.

Amount of under-applied = 380,000 - 377,000 = $3,000

d. Journal enty to record under-applied overhead

Account Title Debit Credit
Cost of goods sold 3,000
Manufacturing overhead 3,000